2 C6H10+ 17 CO2+ 10H2O

if 45 grams of oxygen reacts with excess cyclohexene , how many molecules of carbon dioxide will be formed


1
Expert's answer
2021-03-22T06:27:31-0400

Molar Mass of CO2 = 44g/mol

Molar Mass of cyclohexane = 84.16

= 84.16 × 2 = 168.32/44 = 3

= 3 × (3molCO2/1mol C6H10) × (6.022×10^23 molecules/ mol CO2)

= 5.4198× 10^24 CO2 molecules